OpenStudy (anonymous):

a manufacturer determines that the numbe of drills it can sell is given by formula d=-4p^2+160p-260, where p is price of drills A. at what price will manufacturer sell the maximum of drills? B.what is max number of drills can be sold. show work please :)

OpenStudy (anonymous):

d' = -4* 2 * p + 160 = first differetiate now we take the 1st one and say that d' equals zero 0 = -8p +160 -160 = -8p p = 20 into d(p) d = -4 * 20^2 +160 * 20 - 260 d = -1600 + 320 - 260 d = -1540 At the price of 1540 i would say if im not wrong